8 15. A machine dispenses a liquid drug into bottles in such a way that the standard on of the contents is 54 milliliters. A new machine is tested on a sample of 24 containers and the standard deviation for this sample group is 1 At the 500 ound to be 21 milliliters o level of significance, test the claim that the amounts dispensed by the new e ve asmaller standard deviation (use 2 - distribution for analysis). (7 points). lanl of ew Fneland is concerned about the amount of debt being accru
Answer #1
null hypothesis:     Ho:         σ = 54
Alternate hypothesis: Ha: σ < 54
for sample size n:                      = 24
s2 = 441.00
at 0.05 level and (n-1=23) df crtiical value= 13.091
Decision rule:reject Ho if test statistic F in critical region: X2 < 13.091
test stat : 2=(n-1)s22= 3.48

as test statistic is in rejectiion region we reject null hypothesis and conclude that new machine have a smaller standard deviation